Grade 10Mechanics

a stone projected vertically upward with an initial velocity 128 ft/sec moves according to the law S=128t-16t^2 the acceleration at t=3 seconds is

pani
8 Years agoGrade 10

1 Answer

kiran kumar reddy
8 Years ago
V=ds/dt=128-32t
a=dV/dt =0-32
On solving s in this way you get
a=-32
a=-32ft/s² 
I think it is the answer of this question